OpenStudy (anonymous):

Sin (B-A) is equal to _____,when B= 270 degress and A is an acute angle.

OpenStudy (anonymous):

i thought it is 90 degree but its not because of acute angle

OpenStudy (jamesj):

This an odd question, but in short if A is acute then \[0 \leq A < 90\] and hence B-A is between 180 and 270 degrees. In which case sin(B-A) is a negative number.

OpenStudy (anonymous):

IF the answer is negative the answer is -90 degree?

OpenStudy (jamesj):

No, look at the direction of logical deduction in the statement: "Sin (B-A) is equal to _____,when B= 270 degress and A is an acute angle." That is "If B= 270 degress and A is an acute angle, then Sin (B-A) is equal to _____"

OpenStudy (anonymous):

89 degree?

OpenStudy (jamesj):

"If B= 270 degress and A is an acute angle, then Sin (B-A) is equal to A NEGATIVE NUMBER."

OpenStudy (anonymous):

because of acute angle i confuse....acute angle is less than 90 degree but b-a is a negative number...

OpenStudy (jamesj):

False. If A is between 0 and 90 and B = 270, then B - A is not negative Instead, B - A is between 180 and 270

OpenStudy (jamesj):

Draw a diagram in the plane with a unit circle and see where B-A is. It is an angle that always lies in the 3rd quadrant.

OpenStudy (anonymous):

aha 180 degree?

OpenStudy (jamesj):

What question are you answering with "180 degrees"? I have already given you the answer to the question you posed at the beginning. What are you not understanding? I really think you should draw the diagram and see where the angle B-A is. It is in general an angle in the 3rd quadrant, but it is not one number. It is infinitely many angles, all of which lie in the third quadrant.

OpenStudy (jamesj):

So what are examples of B - A on the diagram you just attached?

OpenStudy (anonymous):

my prof not teach us that before im sorry i just do this for my project but i will i see a lot of degree in there between 90 degree and 270 degree

OpenStudy (jamesj):

In the third quadrant only there is a range of 90 degrees: between 180 and 270 degrees. We are ONLY in the third quadrant because A is an acute angle. Hence A is between 0 and 90 degrees. Hence for B = 270 degrees, B - A is between 180 and 270 degrees.

OpenStudy (anonymous):

i see a acute angle in the diagram 60 degree.. =210

OpenStudy (jamesj):

For instance, yes.

OpenStudy (anonymous):

what yes im correct?

OpenStudy (jamesj):

210 is an example of an angle B - A when B = 270 and A is acute.

OpenStudy (jamesj):

and sin(210) is a negative number.

OpenStudy (jamesj):

In fact sin(270-A) is negative for all acute angles A.

OpenStudy (anonymous):

i hate my professor y did he ask for a project like this with 13 question to go then pass tomorrrow then i feel like im home study i never learn from him

OpenStudy (jamesj):

So what is the answer to the question: Sin (B-A) is equal to _____,when B= 270 degress and A is an acute angle

OpenStudy (anonymous):

-0.5?

OpenStudy (jamesj):

omg. You're not following me at all. Let's go through this again.

OpenStudy (jamesj):

If A is acute it is ANY angle between 0 and 90 degrees. Agreed?

OpenStudy (anonymous):

yes 2x

OpenStudy (jamesj):

For example A = 30 degrees or A = 45 degrees or A = 60 degrees or A= 89.99 degrees or A = 1 degree or A = 25.666858458857564648696745 degrees

OpenStudy (jamesj):

In every case, when A is between 0 and 90 degrees, 270 - A is an angle between 180 and 270 degrees

OpenStudy (jamesj):

and that is the most general thing we can say about that angle.

OpenStudy (jamesj):

When B = 270 degrees and A is acute, B-A can be ANY angle whatsoever between 180 and 270

OpenStudy (jamesj):

That is, B - A is an arbitrary angle in the third quadrant.

OpenStudy (jamesj):

So far, so good?

OpenStudy (anonymous):

i confuse because the answer is negative u said 270-A is an angle between 180 and 270 degrees so far so good the answer is =-210........

OpenStudy (jamesj):

Wait.

OpenStudy (jamesj):

What do we know about the sin of an angle between 180 and 270 degrees?

OpenStudy (jamesj):

What do we know about the sin of an angle in the third quadrant?

OpenStudy (jamesj):

It is always negative.

OpenStudy (anonymous):

yes i see in the picture there all negative

OpenStudy (jamesj):

Hence sin(B-A) is negative when B = 270 and A is acute.

OpenStudy (jamesj):

Therefore the answer to the question: "Sin (B-A) is equal to _____,when B= 270 degress and A is an acute angle" is Sin (B-A) is equal to A NEGATIVE NUMBER,when B= 270 degress and A is an acute angle

OpenStudy (anonymous):

now i know my weakness i cant explain but i know the answer

OpenStudy (anonymous):

sir i have 13 question to go i answer the others but no.13 -15 can u answer TT my head is aching i have 2 hours left to arrange my project TT

OpenStudy (jamesj):

no, I have to go. Do yourself a favor and go back over this problem and make sure you understand it. If you do, I bet you the other questions will be easier.

OpenStudy (anonymous):

ok thank u